# Disentangling Exploration of Large Language Models by Optimal Exploitation

### v1

@@ 4 Experimental Evaluation
In the following section, we describe our experimental setup and results. The experiments aim to demonstrate the utility of our evaluation framework by addressing the following research questions: 1.) Can LLMs sufficiently explore and reduce their exploration gap? and 2.) How do agent instructions impact the exploration and exploitation gaps?
@@ 4.1 Other Statistics
Besides our optimal exploitation framework, defined in section 3.1, we describe exploration using the following statistics:
@@
Agent Return    represents the total return collected by the exploration agent during its interactions with the environment. Although it is not a direct indicator of effective exploration, this metric reflects the agent’s ability to engage in ways that yield cumulative benefits. A higher agent return indicates that the agent is navigating towards rewarding states and is possibly close to exploitation. Contrary to our decomposition, it does not necessarily imply that the agent is efficiently exploring.
@@
State-Space Coverage    reflects the agent’s ability to discover new information and avoid being trapped in local optima. In our task, effective exploration requires covering the entire state space, as balls associated with high rewards are spread randomly on the grid. However, complete exploration may not always be necessary or infeasible in large, infinite, or continuous spaces due to natural limitations and resource constraints. In contrast to state-space coverage, our decomposition also covers cases in which full exploration is not essential.
@@
Memory Redundancy    measures the percentage of overlapping state-action pairs in the memory. The objective should be to minimize unnecessary repetition. In our deterministic environments, the agent should avoid revisiting states or actions as they do not provide new information. Reducing redundancy conserves resources and ensures that the history can be utilized efficiently. Otherwise, information has to be retrieved from longer histories to find novel or informative states to explore.
@@
Sample Efficiency    evaluates the number of interactions required for the agent to converge to its 90% equilibrium of the maximum exploitation return. An efficient agent leverages patterns in the reward and transition dynamics to reach high-reward states as quickly as possible. In our environments, this means minimizing memory redundancy and adding the trajectories to the history such that the exploitation return increase is maximized. Note that the sample efficiency is decoupled from the quality of the final policy. An agent constantly stuck with the same low reward balls is equally valued as another LLM converging to the maximum exploitation return quickly.
4.2 Setup
Our experimental setup [^1] You can find the code under coming soon and test the exploration and exploitation gaps on your own prompts on coming soon directs agents to explore in order to solve tasks. The specific instruction tells to ”collect information that helps to become better at maximizing the return”. We reference this prompt as task-oriented exploration. The exact template is detailed in Appendix 1. Each agent is equipped with a simple memory system that stores the history of past interactions. Notably, unlike many prior studies, our approach does not provide agents with explicit hints about the domain. For the evaluation, we test each open-source model with ten repeated runs and each closed model with five repeated runs on every environment. This results in a total of 30 runs per data point for open-source models and 15 runs for closed models. In the following figures, we report statistical significance using the standard error. Our return decomposition is measured at the end of exploration and as the mean over all episodes. All other statistics are evaluated after all environment interactions. Comprehensive results for all models are included in Appendix B.
To investigate exploration behavior (1), we evaluate agents in Dark Treasure Rooms of varying sizes. To assess the impact of agent instructions (2), we compare our prompt, outlined above, with alternative instructions by incorporating soft lower and upper bounds for the exploration capability of each model. These alternative prompts direct the agent to strictly exploit or aimlessly explore without referencing the task. The instruction impact is analyzed using 7×7 grid environments, with detailed prompt specifications available in Appendix B.5.
@@ 4.3 Models
We evaluate and compare the exploration behavior of the popular open-source LLMs Mistral (7B) (Jiang et al., 2023), Gemma2 (9B and 27B) (Team et al., 2024b), Llama 3.1 (7B) (Dubey et al., 2024), Llama3.3 (70B), Mistral Small (22B) (AI, 2024a), and Mistral Large 2 (123B) (AI, 2024b). We further include closed-models GPT-4o (Achiam et al., 2023), o1-mini (OpenAI, 2024), Gemini 1.5 Pro (Team et al., 2024a), and Claude 3.5 Sonnet (Anthrophic, 2024). We compare the LLMs with a random baseline which we consider the alternative for exploration without prior information.
4.4 Can Large Language Models sufficiently explore?
(a) Agent Return
(b) Exploration Gap
(c) Grid coverage
Figure 2: Comparison of the agent return, exploration gap, and grid coverage. Our exploration gap (see subfigure 2(b), lower is better) measures progress in exploration more accurately than the agent return in 2(a). Note, for example, that Llama3.1 (7B) performs better exploration while achieving lower agent return than Gemma 2 (9B). Also Mistral Large (123B), Llama 3.3 (70B), and o1-mini do not show significant differences in the agent return although they are significantly different (p $\leq 0.01$) in their exploration gap. Our experiments further demonstrate that, generally, coverage after exploration shrinks with grid size (increased long horizon planning).
Most agents fail to minimize their exploration gap and explore the entire state-space.    Our results indicate that LLMs generally struggle to systematically explore the environment, even when their sole objective is to collect information rather than exploit it. The results, averaged across all grid sizes, are presented in table 1. In their final episodes, most models demonstrated significantly larger or indifferent exploration gaps compared to a random baseline, with the exceptions of Mistral Large, o1-mini, and Gemini 1.5 Pro. However, even these models failed to achieve complete state-space coverage, exhibiting non-zero exploration gaps and exploring less than 90% of the state space on average. At closer inspection, all models showed redundant exploration of previously visited states, with exploration disproportionately focused on initial high-reward regions.
@@ 4.4 Can Large Language Models sufficiently explore?
Table 1: Total return gap decomposition averaged over $4\times 4$, $5\times 5$, and $7\times 7$ grids. The exploitation and exploration gaps are normalized to account for varying maximum rewards in different environments. The numbers in parentheses show the fraction that exploitation and exploration gaps make up of the total return gap, respectively.
@@
LLM-based agents are limited in targeted long-horizon exploration.    Our evaluation reveals a statistically significant decrease (p $\leq$ 0.01) in the fraction of state space explored by most models as the grid gets larger from $4\times 4$ to $7\times 7$. Figure 2(c) illustrates the explored state-space coverage for various sizes. Notably, Claude 3.5 Sonnet was the only model to maintain high coverage as the environment size increased, with no significant drop observed (p $\geq$ 0.05). Larger parameter models, including GPT-4o, Mistral Large, and Claude 3.5 Sonnet, also maintained consistent coverage when the grid size increased modestly to $5\times 5$ (p $\geq$ 0.05). Despite achieving superior results in smaller environments and demonstrating strong average state-space coverage, the test-time reasoning model o1-mini similarly experienced a significant performance drop in larger grids.
@@
Weak exploration is not enough.    Contrary to claims in prior work (Huang et al., 2024), our findings demonstrate that smaller LLMs with fewer parameters are unable to sufficiently explore the environment as exploration complexity increases. Our results reveal a clear trend of improved exploration capabilities with model scaling, as evidenced by the statistically significant slope of a linear regression (p $\leq$ 0.01). Interestingly, closed-source models do not necessarily outperform open-source counterparts. For instance, GPT-4o and Llama 3.3 exhibit comparable exploration gaps at the end of training.
@@
Decomposition into optimal gaps measures exploration progress more accurately.    Our results support the proposed decomposition of the total return gap into its optimal exploitation and exploration components. Figures 2(a) and 2(b) illustrate the agent returns and corresponding exploration gaps, respectively. Notably, we find that the agent return often does not correlate with true exploration progress. For instance, models such as Mistral Large (123B), Llama 3.3 (70B), and o1-mini exhibit no significant differences in their agent returns, despite showing statistically significant differences in their exploration gaps (p $\leq$ 0.01). To further investigate this disparity, we conducted an ablation study by substituting our optimal exploration with an alternative LLM. As detailed in Appendix B.1, these results demonstrate that the exploitation return cannot be reliably measured in our environment. These findings underscore our hypothesis that exploitation and exploration are equally challenging and should be assessed independently to ensure an accurate evaluation.
@@
Table 2: Other statistics averaged over $4\times 4$, $5\times 5$, and $7\times 7$ grids. Some LLMs achieve lower agent returns but are higher in their optimal exploitation return (see Gemini 1.5 Pro and Claude 3.5 Sonnet). Generally, we find that all LLM-based agents converge quickly to their individual maximum optimal exploitation return, although more rewarding balls are not explored yet.
4.5 What impact do agent instructions have on exploration?
Exploration stays hard for low parameter models even when the task is ignored.    Models with $\leq$ 9 billion parameters maintain significant exploration gaps even when explicitly prompted to disregard the task and focus solely on exploring unvisited states. For instance, Mistral (7B) and Gemma 2 (9B) exhibit exploration gaps of 57% and 26%, respectively, by the end of exploration. These findings suggest that smaller models face intrinsic limitations in their ability to systematically explore, even under conditions designed to prioritize exploration exclusively. Interestingly, also o1-mini achieves only 82.05% coverage. See appendix B.5.1 for the full results.
Prompts can heavily impact reward collection during exploration.    Our results indicate that prompts can substantially impact reward collection strategies during exploration. Across all LLMs, the exploitation gap increased significantly when agents were prompted with the soft upper bound instruction. Notably, while Mistral Large maintained a similar soft upper bound exploration gap, as shown in figure 3(b), the manner in which this gap was minimized appeared to differ in terms of agent reward dynamics. At closer inspection, the upper bound instruction led to more independence across episodes, whereas task-oriented exploration tended to focus on previously encountered regions with known rewards. These findings highlight the impact of prompt design on exploration and return.
LLMs interpret agent instructions for exploration differently.    Our results reveal that LLMs interpret exploration instructions in diverse ways, leading to varying behaviors. Lower parameter models predominantly exhibit behavior closer to strict exploitation when prompted to collect information to solve the task. In contrast, some models perform exploration more in line with the soft upper bound. Notably, o1-mini achieved lower exploration gaps under task-oriented instructions compared to arbitrary strict exploration. However, it remains an open question how effective aimless exploration is in environments governed by rules or requiring semantic understanding.
(a) Mistral Small (22B)
(b) Mistral Large (123B)
@@ 4.5 What impact do agent instructions have on exploration?
Figure 3: The mean total return gaps decomposed into their exploration and exploitation gaps. larger LLMs employing task-oriented exploration tend to sacrifice more rewards compared to smaller models. Furthermore, LLMs interpret our exploration instruction differently: for Mistral Small, the gap proportions align more closely with the soft lower bound, whereas for Mistral Large, they are closer to the upper bound. These findings support our hypothesis that the proposed evaluation framework can be utilized to analyze agent instructions during prompt engineering.
@@
Our decomposition can be used to gain insights during prompt-construction.    The results demonstrate that our evaluation can effectively analyze various prompting strategies within simulated environments, offering a valuable tool for refining agent behavior prior to deployment in real-world scenarios. This approach represents an initial step toward methods that assess the impact of prompt engineering on the exploration capabilities of LLMs in sequential decision-making.
5 Discussion of Results and Future Implications
@@
Consistent with prior findings (Krishnamurthy et al., 2024; Nie et al., 2024), we observe that most current LLMs struggle with sufficient state-space exploration, even when efforts are made to isolate exploration from exploitation. Contrary to earlier claims (Huang et al., 2024), our results indicate that low-parameter open-source models exhibit significant weaknesses in exploration. This discrepancy may arise from the increased complexity of exploration in our settings compared to simpler household domains. We propose that implementing optimal exploitation evaluations and pursuing advancements in the following directions could represent promising avenues for future research:
@@
Reasoning strategies.    Various reasoning and self-improvement strategies have been proposed recently. Examples, include Reflection (Shinn et al., 2024) and Tree-of-Thoughts (Yao et al., 2024). Evaluating how advanced reasoning approaches influence exploration may be a promising research direction. Our results indicate that test-time reasoning might improve exploration. However, developing strategies specifically designed with exploration in mind may be necessary.
@@
Training for intelligent exploration.    We believe training an agent not only to always exploit but to systematically gather information may be necessary to develop foundational exploration agents. In this regard, LLMs should be explored from a meta-reinforcement learning perspective. Nie et al. (2024) show first evidence that finetuning an agent on exploration trajectories can help in multi-armed bandits. Further, our results demonstrated that exploration capabilities may vary with prompt-engineering. Future agents should be less dependent on prompting and implicitly self-explore.
Memory architectures.    Many agent systems currently focus on developing memory architectures (Wang et al., 2024; Anokhin et al., 2024) that load the LLM context dynamically. Testing such architectures for systematic exploration, similarly to Lu et al. (2024), should be an important dimension in the evaluation. Further, building a meta memory that allows to reuse exploration-specific information across tasks and environments may be a fruitful path.
@@
Small exploration models for real time applications.    We believe giving low parameter models the ability to explore is a necessary steps toward intelligent embedded and real time agents. As larger LLMs have shown to be more proficient at exploration, distillation methods could possibly be developed to teach exploration to smaller models, running on more devices.
Improving sequence understanding. During our experiments, we found that agents increasingly produce invalid actions and misunderstand longer sequences of past actions. Developing agents specifically made for long-sequence understanding is crucial for future LLM-based agents. It would be interesting to see if LLMs finetuned on sequential data (Kim et al., 2024; Zeng et al., 2023; Song et al., 2024) also exhibit better exploration capabilities.

### v2

@@ 4 Experimental Evaluation
Subsequently, we describe our experimental setup and results. The experiments aim to demonstrate the utility of our framework by addressing the following research questions: 1.) Effectiveness: Can LLMs explore and reduce the exploration gap? and 2.) Prompt-engineering: How do agent instructions impact exploration and exploitation gaps?
4.1 Setup
Agent instruction. Our experimental setup directs agents to explore in order to solve tasks. The specific instruction tells to ”collect information that helps to become better at maximizing the return”. We reference this prompt as task-oriented exploration. The exact template is detailed in appendix A.4. Each agent is equipped with a simple memory system that stores the history of past interactions. Notably, unlike many prior studies, our approach does not provide agents with explicit hints about the domain.
Statistical robustness. In terms of statistical robustness, we test each open-source with ten and each closed model with five runs on every environment. This results in a total of 30 runs per data point for open-source models and 15 runs for closed models. In the following Figures, we report statistical significance using the standard error. Our return decomposition is measured at the end of exploration and as the mean over all episodes. The values are normalized to account for environments with different maximum returns. Other statistics are evaluated after all environment interactions. Full results are listed in appendix B and C.
Evaluation procedure. The 1.) Effectiveness is evaluated across three grids of Dark Treasure Rooms ($4\times 4$, $5\times 5$, and $\times 7$) and mazes. The configurations test the scalability of the planning horizon. To assess 2.) Prompt-engineering, we compare our task-oriented agent instruction, with alternative prompts in a 7$\times$7 grid. We set soft lower and upper bounds for the exploration capabilities by directing the agents to strictly exploit or aimlessly explore without referencing the task. See appendix D for more information and the alternative instructions.
@@ 4.2 Other Statistics
We also describe exploration using the following statistics:
@@
Agent return represents the total return collected by the agent during its interactions with the environment. Although it is not a direct indicator of effective exploration, this metric reflects the ability to engage in ways that yield cumulative benefits. A higher agent return implies that the agent is navigating towards rewarding states and is possibly close to exploitation. Contrary to our decomposition, it does not necessarily imply that the agent is effectively exploring.
@@
State-space coverage reflects the ability to discover new information and avoid being trapped in local optima. In our task, effective exploration requires covering the entire state space, as balls with high rewards are randomly spread. However, complete exploration may not always be necessary or could be infeasible in large, infinite, or continuous spaces due to natural and resource constraints. In contrast, our decomposition also covers cases in which full exploration is not essential.
@@
Memory redundancy measures the fraction of redundant state-action pairs in the memory. The objective should be to minimize repetition. In our environments, the agent should avoid revisiting states or actions as they do not provide new information. Reducing redundancy conserves resources and ensures that the history can be utilized efficiently. Otherwise, information has to be retrieved from longer histories to find novel or informative states to explore.
@@
Sample efficiency evaluates the number of environment interactions required to converge to the 90% equilibrium of the maximum achieved exploitation return. This means minimizing memory redundancy and adding trajectories to the history such that the exploitation return increase is maximized. Note that sample efficiency is decoupled from the quality of the final policy. Being constantly stuck with the same low reward balls is equally as fast convergence to the maximum exploitation return.
@@ 4.3 Models
We evaluate the exploration ability of the popular open-source LLMs Mistral (7B) Jiang et al. (2023), Gemma2 (9B and 27B) Team et al. (2024b), Llama 3.1 (7B) Dubey et al. (2024), Llama3.3 (70B), Mistral Small (22B) AI (2024a), and Mistral Large 2 (123B) AI (2024b). We further include closed-models GPT-4o Achiam et al. (2023), o1-mini OpenAI (2024), Gemini 1.5 Pro Team et al. (2024a), and Claude 3.5 Sonnet Anthrophic (2024). The LLMs are compared to a random baseline.
4.4 Can Large Language Models Explore?
Most agents fail to minimize their exploration gap. Our results indicate that LLMs struggle to explore the state-space although their sole objective is to collect information rather than exploit it. The results, averaged across all grid sizes of Dark Treasure Rooms, are presented in Table 1 and 2. In this case, almost all models demonstrated significantly larger or indifferent exploration gaps in their final episodes compared to a random baseline, with the exceptions of Mistral Large, o1-mini, and Gemini 1.5 Pro. However, also these models failed to achieve complete state-space coverage, exhibiting non-zero exploration gaps and exploring less than 90% of the state space on average. Performance was additionally worse in the maze environment, where the increased structural complexity and sparse reward distribution led to lower coverage and larger exploration gaps, further emphasizing the models’ limitations. At closer inspection, models showed redundant exploration of previously visited states, with exploration disproportionately focused on initial high-reward regions.
Figure 2: The agent return, exploration gap, and grid coverage in $5\times 5$ Dark Treasure Rooms. Our exploration gap (middle, lower is better) measures progress in exploration accurately in contrast to the agent return. Note, for example, that Llama3.1 (7B) performs better exploration while achieving lower agent return than Gemma 2 (9B). Also Mistral Large (123B), Llama 3.3 (70B), and o1-mini do not show significant differences in the agent return although they are significantly different (p $\leq 0.01$) in their exploration gap. Our experiments further demonstrate that, coverage shrinks with grid size and, therefore, increased long horizon planning.
Table 1: Total return gap decomposition averaged over the $4\times 4$, $5\times 5$, and $7\times 7$ Dark Treasure Rooms. The exploitation and exploration gaps are normalized to account for varying maximum rewards in different environments. The numbers in parentheses show the fraction that exploitation and exploration gaps make up of the total return gap, respectively.
@@ 4.4 Can Large Language Models Explore?
LLMs are limited in targeted long-horizon exploration. Our evaluation reveals a statistically significant decrease (p $\leq$ 0.01) in the fraction of the explored state-space by most models as the grid gets larger. Figure 2 contains a plot showing the explored state-space coverage for the different sizes grid sizes. Notably, Claude 3.5 Sonnet was the only model to maintain high coverage as the environment size increased, with no significant drop observed (p $\geq$ 0.05). Larger parameter models, including GPT-4o, Mistral Large, and Claude 3.5 Sonnet, only maintained consistent coverage when the grid size increased slightly from $4\times 4$ to $5\times 5$ (p $\geq$ 0.05). Despite achieving superior results and demonstrating strong average state-space coverage in smaller environments, the test-time reasoning model o1-mini similarly experienced a significant performance drop in larger grids. Likewise, all models exhibited the largest exploration gaps in the maze environments, underscoring the limitations of careful long-horizon planning in more complex settings.
@@
Weak exploration is not enough. Contrary to claims in prior work Huang et al. (2024), our findings demonstrate that smaller LLMs with fewer parameters are unable to sufficiently explore the environment. This finding is probably due to the increased exploration complexity in our setting. Our results reveal a clear trend of improved exploration capabilities with model scaling, as evidenced by the statistically significant slope of a linear regression (p $\leq$ 0.01). Interestingly, closed-source models do not necessarily outperform open-source counterparts. For instance, GPT-4o and Llama 3.3 exhibit comparable exploration gaps at the end of training.
@@
Optimality gaps measure exploration accurately. Our results support the decomposition of the total return gap into its optimal exploitation and exploration components. Reconsider in this context the agent returns and the corresponding exploration gaps, which are illustrated in Figure 2. We find that the agent return does often not correlate with true exploration progress. For instance, models such as Mistral Large (123B), Llama 3.3 (70B), and o1-mini exhibit no significant differences in their agent returns, despite showing statistically significant differences in their exploration gaps (p $\leq$ 0.01). To further investigate this disparity, we conducted an ablation study in which we substituted our optimal exploration with an alternative LLM. The results for this follow-up experiment are listed in appendix A.7. The additional ablation demonstrates that the exploitation return cannot be reliably measured in this setting. These findings underscore our hypothesis that exploitation and exploration are equally challenging and should be assessed independently to ensure an accurate evaluation.
4.5 What is the Impact of Agent Instructions?
Exploration stays hard for low parameter models. In the following, we compare our task-oriented exploration prompt with the alternative instructions as described in Section 4.1. The results show that models with $\leq$ 9 billion parameters still maintain significant exploration gaps when prompted to disregard the task and focus solely on exploring any unvisited states. For instance, Mistral (7B) and Gemma 2 (9B) exhibit exploration gaps of 57% and 26%, respectively, in their last episode. These findings may suggest that smaller models face intrinsic limitations in their ability to explore, even when tasked to prioritize exploration exclusively. Interestingly, also more competent models, such as o1-mini, do not achieve full coverage. The full list of results for the soft upper bound prompt are detailed in appendix D.2. The soft lower bound outcomes can be found in appendix D.1.
Table 2: The other statistics averaged over the $4\times 4$, $5\times 5$, and $7\times 7$ Dark Treasure Rooms. Some LLMs achieve lower agent returns but are higher in their optimal exploitation return (see Gemini 1.5 Pro and Claude 3.5 Sonnet). We find that all agents converge quickly to their individual maximum optimal exploitation return, although more rewarding balls are not explored yet.
@@ 4.5 What is the Impact of Agent Instructions?
Figure 3: The normalized mean total return gaps decomposed into their exploration and exploitation components. Larger LLMs employing task-oriented exploration tend to sacrifice more rewards compared to smaller models. Furthermore, LLMs interpret our exploration instruction differently: for Mistral Small, the gap proportions align more closely with the soft lower bound, whereas for Mistral Large, they are closer to the upper bound. These findings support our hypothesis that our evaluation can be utilized to analyze agent instructions during prompt engineering.
Prompts can heavily impact reward collection. Our results indicate that prompts can impact reward collection during exploration. The exploitation gap increased significantly when agents were prompted with the soft upper bound instruction that is tailored towards strictly performing any exploration. Figure 3 contains the decomposed total return gaps of Mistral Large. While the model maintained a similar task-oriented exploration gap, the manner in which this gap was minimized appeared to differ in terms of agent returns. At closer inspection, the upper bound instruction leads to more independence across episodes, whereas task-oriented exploration tends to focus on previously visited regions. These findings highlight the impact of prompt design on exploration and return.
LLMs interpret the agent instructions differently. Our results suggest that LLMs interpret the same task-oriented exploration instruction with varying behaviors. Lower parameter models predominantly exhibit behavior closer to strict exploitation when prompted to collect information to solve the task. In contrast, other models perform task-oriented exploration more in line with the soft upper bound. Surprisingly, o1-mini achieved lower exploration gaps under task-oriented instructions compared to arbitrary strict exploration. However, it remains an open question how effective aimless exploration is in environments governed by rules or requiring semantic understanding.
@@
Our framework assists prompt-engineering. The results demonstrate that our evaluation can effectively analyze various prompting strategies within simulated environments. We therefore believe it is a valuable tool for refining agent behavior prior to deployment in real-world scenarios. This approach represents an initial step toward methods that assess the impact of prompt engineering on the exploration capabilities of LLMs in sequential decision-making.
5 Discussion and Future Implications
@@
Consistent with prior findings Krishnamurthy et al. (2024); Nie et al. (2024), we observe that most of the popular LLMs struggle with state-space exploration when exploration is isolated from exploitation. Contrary to earlier claims Huang et al. (2024), our results indicate that low parameter models exhibit significant weaknesses. This discrepancy may arise from the increased complexity to find objects in our setting while the small number of rooms is easily covered in household domains. Moreover, our experiments have shown that comparing models based on agent returns or an LLM for the exploitation is not reliable. Our evaluation procedure fills this gap and allows a fair evaluation of the following promising avenues for future research:
@@
Reasoning and self-improvement. Various reasoning and self-improvement strategies have been proposed recently. Examples include Reflection Shinn et al. (2024), Tree-of-Thoughts Yao et al. (2024), and RL Guo et al. (2025). Evaluating how these approaches influence exploration may be a promising research direction. Our results indicate that test-time reasoning has a positive impact. In this regard, it would also be interesting to see, if LLMs can self-improve their own exploration skills. Developing strategies specifically designed with exploration in mind could be necessary.
@@
Training for intelligent exploration. Training an agent to systematically gather information may be required to develop foundational agents. Hereby, LLMs should be investigated from a meta-RL perspective. Nie et al. (2024) show first evidence that finetuning an agent on exploration trajectories can help in multi-armed bandits. Our results demonstrate that exploration capabilities vary with prompt-engineering. Future agents should be less dependent on prompting and implicitly explore.
@@
Small exploration models for real time applications.    Our results have revealed new research directions for low parameter LLMs. We believe improving the ability of weak models to explore is a necessary steps towards many real life applications, requiring embedded and real time agents. As larger LLMs have shown to be more proficient at exploration, distillation methods could possibly be developed to teach exploration to smaller models.
We discuss additional future implications in appendix E.